Find the length of an air filled tube closed on one end and open on the other if its fundamental frequency is 240 Hz
Answer #1

Fundamental frequency of closed pipe

f=v/4l

v=velocity of sound =343m/s

l=length of pipe =?

f=240Hz

f=v/4l

240=343/4l

l=343/(4X240)=0.357meter =35.7cm